OpenStudy (anonymous):

What is the distance between the points (-5, 2) and (4, 9)? .

OpenStudy (sriram):

sqrt(130)

OpenStudy (anonymous):

\[\sqrt{(9-2)^2+(4-(-5)^2}=\sqrt{7^2+9^2}\]

OpenStudy (sriram):

distance =sqrt( {-5-4}^2 + {2-9}^2 ) =sqrt( 7^2 + 9^2 ) =sqrt(130)

OpenStudy (anonymous):

\[\sqrt{49+81}=\sqrt{130}\]

OpenStudy (anonymous):

\[=10\sqrt{3}\]

OpenStudy (anonymous):

oh hold the phone

OpenStudy (anonymous):

oh hold the phone

OpenStudy (anonymous):

\[\sqrt{130}\] is right, but \[10\sqrt{3}\] is not \[\sqrt{300}=10\sqrt{3}\]

OpenStudy (anonymous):

utilize the distance formul, and plug in your points